I NEED HELP COULD ANYONE SHOW ME HOW THESE ARE DONE? QUESTION READS FIND THE EQUATION OF THE LINE DETERMINED BY THE GIVEN INFORMATION. 1. M=1; P=(-7,0) 2. P1=(4,-5); P2=(11,2) 3. P1=(0,6); P2=(-2,3) 4. M=2; P=(0,1)
Answer by stanbon(75887)

FIND THE EQUATION OF THE LINE DETERMINED BY THE GIVEN INFORMATION. 1. M=1; P=(-7,0) 2. P1=(4,-5); P2=(11,2) 3. P1=(0,6); P2=(-2,3) 4. M=2; P=(0,1)
-----------------
Line equations look like this: y = mx + b
m is the slope
b is the y-intercept; it is the value of y when x = 0.
----------------
1. M=1; P=(-7,0)
You are told the slope is 1 and y = 0 when x = -7
Put those numbers into y = mx + b to solve for "b":
0 = 1*-7 + b
b = 7
EQUATION: y = x + 7
------------------------
2. P1=(4,-5); P2=(11,2) 3. P1=(0,6); P2=(-2,3)
These cannot be worked because you only have one point to work with
------------------------
4. M=2; P=(0,1)
slope = 2
1 = 2*0 + b
b = 1
EQUATION:
y = 2x + 1
=================
Cheers,
Stan H.